How much should be invested today in order to have a $18,000 balance 3 years from now if the investments pays 9% annual interest with interest compounded monthly?
Answer by Boreal(15235)

18000=Po(1+(.09/12))^36, the 36 being the total number of compoundings.
18000=Po(1.308) but don't round until end.
Po=18000/(1.30864...)=$13754.68